OpenStudy (prince1342):

If f(x) = 2x2 + 2, find f(5). 102 52 22 -1.5

undeadknight26 (undeadknight26):

okay so im assuming its 2x^2?

OpenStudy (prince1342):

yep

undeadknight26 (undeadknight26):

okay so simply plug in 5 for x and solve.

OpenStudy (prince1342):

wut

undeadknight26 (undeadknight26):

f(x) = 2x^2 + 2 f(5) = 2 x 5^2 + 2 f(5) = 2 x 25 + 2 f(5) = 50 + 2 f(5) = 52

OpenStudy (prince1342):

THANK YOUUU

OpenStudy (prince1342):

If f(x) is the height, in cm, of a sunflower plant which is x days old, which of the following statements best describes the meaning of f(60) = 210? The height of the sunflower plant is 60 cm when it is 210 days old. The height of the sunflower plant is 210 cm when it is 60 days old. The height of the sunflower plant is 210 cm when it is 3.5 days old. The height of the sunflower plant is 60 cm when it is 3.5 days old.

undeadknight26 (undeadknight26):

No problemo.

undeadknight26 (undeadknight26):

quite sure its A but not 100%.

OpenStudy (prince1342):

The first five terms of a sequence are shown below. 5, 9, 13, 17, 21 If the nth term of this sequence is represented by f(n), which of the following functions best represents this sequence? f(n) = 5n - 1; n ≥ 1 f(n) = 5n + 1; n ≥ 1 f(n) = 4n - 1; n ≥ 1 f(n) = 4n + 1; n ≥ 1